# Arithmetical enhancements of the Kogbetliantz method for the SVD of order two

### v2

@@ 1 Introduction
The singular value decomposition (SVD) of a square matrix of order two is a widely used numerical tool. In LAPACK [1] alone, its xLASV2 routine for the SVD of real upper triangular $2\times 2$ matrices is a building block for the QZ algorithm [2] for the generalized eigenvalue problem $Ax=\lambda Bx$, with $A$ and $B$ real and square, and for the SVD of real bidiagonal matrices by the implicit QR algorithm [3]. Also, the oldest method for the SVD of square matrices that is still in use was developed by Kogbetliantz [4], based on the SVD of order two, and as such is the primary motivation for this research.
@@ 2 Floating-point considerations
[…]
@@
Radix two is assumed for floating-point values. Scaling of a value $x$ by $2^{s}$ where $s\in\mathbb{Z}$ is exact if the result is normal. Only non-normal results can lose precision. Let, for $x=\pm 0$, $e_{x}=0$ and $f_{x}=0$, and for a finite non-zero $x$ let the exponent be $e_{x}=\lfloor\lg|x|\rfloor$ and the “mantissa” $21\leq|f_{x}|<$ 2, such that $x=\mathop{\mathrm{fl}}(2^{e_{x}}f_{x})$. Also, let $f_{x}=x$ for $x=\pm\infty$, while $e_{x}=0$. Note that $f_{x}$ is normal even for subnormal $x$. Keep in mind that the frexp routine represents a finite non-zero $x$ with $e_{x}^{\prime}=e_{x}+1$ and $f_{x}^{\prime}=f_{x}/2$.
Let $\mu$ be the smallest and $\nu$ the largest positive finite normal value. Then, in the notation just introduced, $e_{\mu}=\lg\mu=-126$ or $-1022$, and $e_{\nu}=\lfloor\lg\nu\rfloor=127$ or $1023$, for single or double precision. Lemma 2.1 can now be stated using this notation.
@@
For the second equality in (2), note that $(\nu+1)^{2}=\nu^{2}+1+2>\nu^{2}+1>\nu^{2}$ since $\nu>0$. By taking the square roots, it follows that $\nu+1>\sqrt{\nu^{2}+1}>\nu$, and therefore
@@
An underlined expression denotes a computed floating-point approximation of the exact value of that expression. Assuming $40\leq|\phi|\leq\pi/$ 4, and with $\tan(2\phi)$ and $\underline{\tan(2\phi)}$ given, where the latter may or may not be equal to $\mathop{\mathrm{fl}}(\tan(2\phi))$, $\tan\phi$ and $\underline{\tan\phi}$ are
@@
and the relation $\prec$, that compares the represented values in the $<$ sense, is given as
@@
[…]
@@ 3 The SVD of general matrices of order two
In general, $U$ is a product of permutations (denoted by $P$ with subscripts and including $I$), sign matrices (denoted by $S$ with subscripts) with each diagonal element being either $1$ or $-1$ while the rest are zeros, and plane rotations by the angles $\vartheta$ and $\varphi$. If $U_{\vartheta}$ is not generated, the notation changes from $U_{\varphi}$ to $U_{\phi}$. Likewise, $V$ is a product of permutations, a sign matrix, and a plane rotation by the angle $\psi$, where
@@
[…]
@@ 3.1 Prescaling of the matrix and the simple cases ($\mathfrak{t}\cong 3,5,9$)
This proof and the following one use several techniques from [21, Theorem 1]. Due to the structure of a matrix that $U_{\theta}^{T}$ or $V_{\theta}$ is applied to, containing in each row and column one zero and one ±$\pm 1$, it follows that $U$ or $V$ have in each row and column ±$\pm\underline{\cos\theta}$ and ±$\pm\underline{\sin\theta}$, computed implicitly, for which Lemma 3.2 gives error bounds.
@@
If, in (14), $\underline{\tan\theta}<\mu$, then $\underline{\sec\theta}=1$ since $11\leq\underline{\sec\theta}\leq\mathop{\mathrm{fl}}(\sqrt{1+\mu^{2}})\leq\mathop{\mathrm{fl}}(1+\mu)=$ 1, and the relative error in $\underline{\sec\theta}$ is below $\varepsilon$ for any standard floating-point datatype. Thus, even though $\underline{\tan\theta}$ can be relatively inaccurate, (16) holds for all $\underline{\sec\theta}$. Also, $\underline{\cos\theta}$ is always relatively accurate, but $\underline{\sin\theta}$ might not be if $\underline{\tan\theta}<\mu$, when $\underline{\sin\theta}=\underline{\tan\theta}$.
3.2 A (pivoted) $URV$ factorization of order two ($\mathfrak{t}^{\prime}\cong 13,15$)
@@ 3.2.2 A fully pivoted URV when $\mathfrak{t}^{\prime}=15$
However, $\underline{U_{+}^{T}}$ is not explicitly formed, as explained in Section 3.3. Now $\mathfrak{t}^{\prime\prime}=13$ for $\underline{R}$.
@@
To prove (23) and (24), solve $x\pm y\delta_{\vartheta}\tan\vartheta=(x\pm y\tan\vartheta)(1+\epsilon_{\pm})$ for ϵ±subscriptitalic-$\epsilon_{\pm}$ with $xy\neq 0$. After expanding and rearranging the terms, it follows that
@@
It remains to prove (34), since the other relations follow from Lemma 3.3.
@@
+¯$\underline{\star_{+}}&=\mathop{\mathrm{FL}}(\star_{+})=\star_{+}(1+\epsilon_{+}^{\prime}),\quad\star_{+}=g_{12}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}+g_{22}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{+}^{\prime}|\leq\varepsilon_{\text{{T}}},\\ \underline{\star_{-}}&=\mathop{\mathrm{FL}}(\star_{-})=\star_{-}(1+\epsilon_{-}^{\prime}),\quad\star_{-}=g_{22}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}-g_{12}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{-}^{\prime}|\leq\varepsilon_{\text{{T}}},\end{aligned}\\ \mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))=\mathop{\mathrm{fl}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})(1+\epsilon_{\pm}^{\prime})(1+\epsilon_{/}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})\delta_{\pm}^{\prime},\end{gathered}$ (35)
with $|\epsilon_{/}^{\prime\prime\prime}|\leq\varepsilon$, since the doubly-rounded quotient $\mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))$ is correctly rounded to the working precision from the value of ±¯$\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}$, where $\star_{\pm}$ can be replaced by $\star_{+}$ or $\star_{-}$. This double rounding is innocuous with the standard datatypes, i.e, when rounding from exact to double to single and from exact to quadruple to double precision, due to [26, Theorem 4] and [27, Theorem 29 and Table II], but might not be for specialized, non-standard datatypes, for which the proposed method would have to be redesigned.
@@
By fixing the sign to $+$ or -- in the $\pm$ subscripts in (35) and (36), let $\delta_{3}^{\prime}=\delta_{-}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$ and $\delta_{3}^{\prime\prime}=\delta_{+}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$. Now bound $\delta_{3}^{\prime}$ and $\delta_{3}^{\prime\prime}$ below by $\delta_{3}^{-}$ and above by $\delta_{3}^{+}$, by combining the appropriate lower and upper bounds for $\delta_{\pm}^{\prime}$, $\delta_{\vartheta}^{\prime}$ from Lemma 3.1, and $\delta_{/}^{\prime\prime\prime\prime}$, where
@@
what proves (34), by minimizing the numerator and maximizing the denominator to minimize the expression for $\delta_{3}^{\prime}$ and $\delta_{3}^{\prime\prime}$, and vice versa, as in the proof of Lemma 3.3. ∎
Therefore, it is possible to compute $\underline{R}$ with high relative accuracy in the absence of underflows, in the working precision only, but it is easier to employ two multiplications, one addition or subtraction, and one division in a wider, more precise datatype. Table 1 shows by how many $\varepsilon$s the lower and the upper bounds for $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ from (23), (24), and (34), respectively, differ from unity. The quantities in the table’s header were computed symbolically as algebraic expressions by substituting $\varepsilon$ and $\varepsilon_{\text{{T}}}$ with their defining powers of two, then approximated numerically with $p_{\text{{T}}}$ decimal digits, and rounded upwards to nine digits after the decimal point, by a Wolfram Language script [^4] The relerr.wls file in the code supplement, executed by the Wolfram Engine 14.0.0 for macOS (Intel)..
@@ 3.3.1 The SVD of $\widetilde{R}$
28:  $\underline{\sec\varphi}=\mathop{\mathrm{hypot}}(\underline{\tan\varphi},1);\quad\underline{\cos\varphi}=\mathop{\mathrm{fl}}(1/\underline{\sec\varphi});\quad\underline{\sin\varphi}=\mathop{\mathrm{fl}}(\underline{\tan\varphi}/\underline{\sec\varphi})$
@@
1:  $\underline{t}=\mathop{\mathrm{fma}}(\underline{\tilde{r}_{22}},\underline{\tan\varphi},\underline{\tilde{r}_{12}});\quad\underline{\tan\psi}=\mathop{\mathrm{fl}}(\underline{t}/\underline{\tilde{r}_{11}})$ // (43)
@@ 3.3.2 The SVD of $G$
For (51), $P_{U}^{T}S_{1}^{T}$ from (21) is explicitly built and stored. It contains exactly one ±$\pm 1$ element in each row and column, while the other is zero. Its multiplication by $\underline{U_{\bm{\varphi}\pm\vartheta}^{T}}$ is thus performed error-free. The tangents computed as in (50) might be relatively inaccurate in theory, but the transformations they define via the cosines and the sines from either (48) or (49) are numerically orthogonal in practice, as shown in Section 4.
@@ 3.4 A complex input matrix
If $G$ is purely imaginary, ±$\pm\mathrm{i}G$ is real. Else, if $G$ has at least one complex element, the proposed real method is altered, as detailed in [14, 15], in the following ways:
@@ 4 Numerical testing
Let the LAPACK’s xLASV2 routine be denoted by $\mathtt{L}$. The Kogbetliantz SVD in the same datatype is denoted by $\mathtt{K}$. Unless such information is clear from the context, let the results’ designators carry a subscript $\mathtt{K}$ or $\mathtt{L}$ in the following figures, depending on the routine that computed them, and also a superscript $\circ$ or $\bullet$, signifying how the input matrices were generated. All inputs were random. Those denoted by $\circ$ had their elements generated as Fortran’s pseudorandom numbers not above unity in magnitude, and those symbolized by $\bullet$ had their elements’ magnitudes in the “safe” range $[\mu,\nu/4]$, as defined by (11), to avoid overflows with $\mathtt{L}$ and underflows due to the prescaling in $\mathtt{K}$. The latter random numbers were provided by the CPU’s rdrand instructions. If not kept, the $\bullet$ inputs are thus not reproducible, unlike the $\circ$ ones if the seed is preserved.
@@
and the departure from orthogonality in the Frobenius norm for matrices of the left and right singular vectors (what can be seen as the relative error with respect to $I$) as
@@
(ignoring the exponent of 00) it might be expected that $\mathtt{K}$ computes $\underline{\sigma_{2}}$ accurately, while $\mathtt{L}$ should additionally be safeguarded by its user from the elements too close to $\mu$. The proposed prescaling, but with $\mathfrak{s}_{\mathtt{L}}=\mathfrak{s}+1$ (or more), might be applied to $G$ before $\mathtt{L}$.
@@
[…]

### v3

@@ 1 Introduction
The singular value decomposition (SVD) of a square matrix of order two is a widely used numerical tool. In LAPACK [1] alone, its xLASV2 routine for the SVD of real upper triangular $2\times 2$ matrices is a building block for the QZ algorithm [2] for the generalized eigenvalue problem $Ax=\lambda Bx$, with AA and $B$ real and square, and for the SVD of real bidiagonal matrices by the implicit QR algorithm [3]. Also, the oldest method for the SVD of square matrices that is still in use was developed by Kogbetliantz [4], based on the SVD of order two, and as such is the primary motivation for this research.
@@ 2 Floating-point considerations
Let $x$ be a real, infinite, or undefined (Not-a-Number) value: $x\in\mathbb{R}\cup\{-\infty,+\infty,\mathrm{NaN}\}$. Its floating-point representation is denoted by $\mathop{\mathrm{fl}}(x)$ and is obtained by rounding $x$ to a value of the chosen standard floating-point datatype using the rounding mode in effect, that is here assumed to be to nearest (ties to even). If the result is normal, $\mathop{\mathrm{fl}}(x)=x(1+\epsilon)$, where $|\epsilon|\leq\varepsilon=2^{-p}$ and $p$ is the number of bits in the significand of a floating-point value. In the LAPACK’s terms, $\varepsilon_{\text{{x}}}=\text{{xLAMCH('E')}}$, where $\text{{x}}=\text{{S}}$ or D. Thus, $p_{\text{{x}}}=24$ or $53$, and $\varepsilon_{\text{{x}}}=2^{-24}$ or $2^{-53}$ for single (S) or double (D) precision, respectively. The gradual underflow mode, allowing subnormal inputs and outputs, has to be enabled (e.g., on Intel-compatible architectures the Denormals-Are-Zero and Flush-To-Zero modes have to be turned off). Trapping on floating-point exceptions has to be disabled (what is the default non-stop handling from [18, Sect. 7.1]).
@@
Radix two is assumed for floating-point values. Scaling of a value $x$ by $2^{s}$ where $s\in\mathbb{Z}$ is exact if the result is normal. Only non-normal results can lose precision. Let, for $x=\pm 0$, $e_{x}=0$ and $f_{x}=0$, and for a finite non-zero $x$ let the exponent be $e_{x}=\lfloor\lg|x|\rfloor$ and the “mantissa” $1\leq|f_{x}|<2$, such that $x=\mathop{\mathrm{fl}}(2^{e_{x}}f_{x})$. Also, let $f_{x}=x$ for $x=\pm\infty$, while $e_{x}=0$. Note that $f_{x}$ is normal even for subnormal $x$. Keep in mind that the frexp routine represents a finite non-zero $x$ with $e_{x}^{\prime}=e_{x}+1$ and $f_{x}^{\prime}=f_{x}/2$.
Let $\mu$ be the smallest and $\nu$ the largest positive finite normal value. Then, in the notation just introduced, $e_{\mu}=\lg\mu=-126$ or 1022-1022, and $e_{\nu}=\lfloor\lg\nu\rfloor=127$ or $1023$, for single or double precision. Lemma 2.1 can now be stated using this notation.
@@
For the second equality in (2), note that $(\nu+1)^{2}=\nu^{2}+1+2\nu>\nu^{2}+1>\nu^{2}$ since $\nu>0$. By taking the square roots, it follows that $\nu+1>\sqrt{\nu^{2}+1}>\nu$, and therefore
@@
An underlined expression denotes a computed floating-point approximation of the exact value of that expression. Assuming $0\leq|\phi|\leq\pi/4$, and with $\tan(2\phi)$ and $\underline{\tan(2\phi)}$ given, where the latter may or may not be equal to $\mathop{\mathrm{fl}}(\tan(2\phi))$, $\tan\phi$ and $\underline{\tan\phi}$ are
@@
and the relation $\prec$, that compares the represented values in the << sense, is given as
@@
$e_{G}=\max_{1\leq i,j\leq n}e_{ij},\quad e_{ij}=\max\{\left\lfloor\lg g_{ij}\right\rfloor,\text{{INT\_MIN}}\}.$ (9)
@@ 3 The SVD of general matrices of order two
In general, $U$ is a product of permutations (denoted by $P$ with subscripts and including II), sign matrices (denoted by $S$ with subscripts) with each diagonal element being either 11 or 1-1 while the rest are zeros, and plane rotations by the angles $\vartheta$ and $\varphi$. If $U_{\vartheta}$ is not generated, the notation changes from $U_{\varphi}$ to $U_{\phi}$. Likewise, $V$ is a product of permutations, a sign matrix, and a plane rotation by the angle $\psi$, where
@@
Depending on its pattern of zeros, a matrix of order two falls into one of the 16 types $\mathfrak{t}$ shown in (13), where $\circ=0$ and $\bullet\neq 0$. Some types are permutationally equivalent to others, what is denoted by $\mathfrak{t}_{1}\cong\mathfrak{t}_{2}$, and means that a $\mathfrak{t}_{1}$-matrix can be pre-multiplied and/or post-multiplied by permutations to be transformed into a $\mathfrak{t}_{2}$-matrix, and vice versa, keeping the number of zeros intact. Each $\mathfrak{t}\neq 0$ has its associated scale type $\mathfrak{s}$.
@@ 3.1 Prescaling of the matrix and the simple cases ($\mathfrak{t}\cong 3,5,9$)
This proof and the following one use several techniques from [21, Theorem 1]. Due to the structure of a matrix that $U_{\theta}^{T}$ or $V_{\theta}$ is applied to, containing in each row and column one zero and one $\pm 1$, it follows that $U$ or $V$ have in each row and column $\pm\underline{\cos\theta}$ and $\pm\underline{\sin\theta}$, computed implicitly, for which Lemma 3.2 gives error bounds.
@@
If, in (14), $\underline{\tan\theta}<\mu$, then $\underline{\sec\theta}=1$ since $1\leq\underline{\sec\theta}\leq\mathop{\mathrm{fl}}(\sqrt{1+\mu^{2}})\leq\mathop{\mathrm{fl}}(1+\mu)=1$, and the relative error in $\underline{\sec\theta}$ is below $\varepsilon$ for any standard floating-point datatype. Thus, even though $\underline{\tan\theta}$ can be relatively inaccurate, (16) holds for all $\underline{\sec\theta}$. Also, $\underline{\cos\theta}$ is always relatively accurate, but $\underline{\sin\theta}$ might not be if $\underline{\tan\theta}<\mu$, when $\underline{\sin\theta}=\underline{\tan\theta}$.
3.2 A (pivoted) URVURV factorization of order two ($\mathfrak{t}^{\prime}\cong 13,15$)
@@ 3.2.2 A fully pivoted URV when $\mathfrak{t}^{\prime}=15$
However, $\underline{U_{+}^{T}}$ is not explicitly formed, as explained in Section $3.$ Now $\mathfrak{t}^{\prime\prime}=13$ for $\underline{R}$.
@@
To prove (23) and (24), solve $x\pm y\delta_{\vartheta}\tan\vartheta=(x\pm y\tan\vartheta)(1+\epsilon_{\pm})$ for $\epsilon_{\pm}$ with $xy\neq 0$. After expanding and rearranging the terms, it follows that
@@
It remains to prove (34), since the other relations follow from Lemma $3.$
@@
$\begin{gathered}\begin{aligned} \underline{\star_{+}}&=\mathop{\mathrm{FL}}(\star_{+})=\star_{+}(1+\epsilon_{+}^{\prime}),\quad\star_{+}=g_{12}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}+g_{22}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{+}^{\prime}|\leq\varepsilon_{\text{{T}}},\\ \underline{\star_{-}}&=\mathop{\mathrm{FL}}(\star_{-})=\star_{-}(1+\epsilon_{-}^{\prime}),\quad\star_{-}=g_{22}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}-g_{12}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{-}^{\prime}|\leq\varepsilon_{\text{{T}}},\end{aligned}\\ \mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))=\mathop{\mathrm{fl}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})(1+\epsilon_{\pm}^{\prime})(1+\epsilon_{/}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})\delta_{\pm}^{\prime},\end{gathered}$ (35)
with $|\epsilon_{/}^{\prime\prime\prime}|\leq\varepsilon$, since the doubly-rounded quotient $\mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))$ is correctly rounded to the working precision from the value of $\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}$, where $\star_{\pm}$ can be replaced by $\star_{+}$ or $\star_{-}$. This double rounding is innocuous with the standard datatypes, i.e, when rounding from exact to double to single and from exact to quadruple to double precision, due to [26, Theorem 4] and [27, Theorem 29 and Table II], but might not be for specialized, non-standard datatypes, for which the proposed method would have to be redesigned.
@@
By fixing the sign to ++ or - in the $\pm$ subscripts in (35) and (36), let $\delta_{3}^{\prime}=\delta_{-}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$ and $\delta_{3}^{\prime\prime}=\delta_{+}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$. Now bound $\delta_{3}^{\prime}$ and $\delta_{3}^{\prime\prime}$ below by $\delta_{3}^{-}$ and above by $\delta_{3}^{+}$, by combining the appropriate lower and upper bounds for $\delta_{\pm}^{\prime}$, $\delta_{\vartheta}^{\prime}$ from Lemma 3.1, and $\delta_{/}^{\prime\prime\prime\prime}$, where
@@
what proves (34), by minimizing the numerator and maximizing the denominator to minimize the expression for $\delta_{3}^{\prime}$ and $\delta_{3}^{\prime\prime}$, and vice versa, as in the proof of Lemma $3.$ ∎
[…]
@@ 3.3.1 The SVD of $\widetilde{R}$
[…]
@@
[…]
@@ 3.3.2 The SVD of $G$
For (51), $P_{U}^{T}S_{1}^{T}$ from (21) is explicitly built and stored. It contains exactly one $\pm 1$ element in each row and column, while the other is zero. Its multiplication by $\underline{U_{\bm{\varphi}\pm\vartheta}^{T}}$ is thus performed error-free. The tangents computed as in (50) might be relatively inaccurate in theory, but the transformations they define via the cosines and the sines from either (48) or (49) are numerically orthogonal in practice, as shown in Section 4.
@@ 3.4 A complex input matrix
If $G$ is purely imaginary, $\pm\mathrm{i}G$ is real. Else, if $G$ has at least one complex element, the proposed real method is altered, as detailed in [14, 15], in the following ways:
@@ 4 Numerical testing
Let the LAPACK’s xLASV2 routine be denoted by $\mathtt{L}$. The Kogbetliantz SVD in the same datatype is denoted by $\mathtt{K}$. Unless such information is clear from the context, let the results’ designators carry a subscript $\mathtt{K}$ or $\mathtt{L}$ in the following figures, depending on the routine that computed them, and also a superscript $\circ$ or $\bullet$, signifying how the input matrices were generated. All inputs were random. Those denoted by $\circ$ had their elements generated as Fortran’s pseudorandom numbers not above unity in magnitude, and those symbolized by $\bullet$ had their elements’ magnitudes in the “safe” range $\mu,\nu/4]$, as defined by (11), to avoid overflows with $\mathtt{L}$ and underflows due to the prescaling in $\mathtt{K}$. The latter random numbers were provided by the CPU’s rdrand instructions. If not kept, the $\bullet$ inputs are thus not reproducible, unlike the $\circ$ ones if the seed is preserved.
@@
and the departure from orthogonality in the Frobenius norm for matrices of the left and right singular vectors (what can be seen as the relative error with respect to II) as
@@
(ignoring the exponent of 0) it might be expected that $\mathtt{K}$ computes $\underline{\sigma_{2}}$ accurately, while $\mathtt{L}$ should additionally be safeguarded by its user from the elements too close to $\mu$. The proposed prescaling, but with $\mathfrak{s}_{\mathtt{L}}=\mathfrak{s}+1$ (or more), might be applied to $G$ before $\mathtt{L}$.
@@
A timing comparison of xLASV2 ($\mathtt{L}$) and the proposed method ($\mathtt{K}$) in single [^6] These results were computed using the Intel oneAPI Base and HPC toolkits, version 2025.0.4. and double precision is given in Table 2. Each method call was timed by the clock_gettime Linux function [^7] See, e.g., https://man7.org/linux/man-pages/man2/clock_gettime.2.html URL. using the CLOCK_MONOTONIC_RAW clock. By construction, $\mathtt{K}$ is more computationally complex than $\mathtt{L}$, so it is not unexpected to be, on average, more than twice slower. In the Kogbetliantz SVD for matrices of order $n>2$, each $2\times 2$ SVD, itself of a constant complexity, is followed by (at most four) $\mathop{O}(n)$ matrix updates, and thus the run-time of any reasonably fast $2\times 2$ SVD method is not a bottleneck. A sufficiently accurate SVD of order two might even reduce somewhat the overall number of iterations of the Kogbetliantz SVD on certain inputs, compared to using a less accurate but faster method, and, if so, partly compensate for its relative slowness.